In the line spectrum of hydrogen there is a ultraviolet line with wavelength 383.5 nm. The frequency of this line is ______ Hz. The energy is _____ J per photon, or ______ kJ/mol. This line corresponds to a transition from the excited state with n=9 to a lower state with n=2. The energy of the lower state is -5.45×10-19 J, so the energy of the excited state with n=9 must be ______ J.
Answer #1

frequency=7.82×10^14 Hz

Energy of one photon=5.18×10^-19J

Energy of one mole photon=312.24KJ/mol

Energy of n=9 state= -10.63×10^-19J